Find the points of intersection of y=2x+4 and y^2=4x+8

Answer by satyareddy22(84)

y=2x+4 ------- (1)
y^2=4x+8= 2(2x+4)-----(2)
but from (1) 2x+4=y
substituting (1) in (2)
y^2=2y
y^2-2y=0
y(y-2)=0
=>y=0,y=2
now substituting y values in (1)
x=-2,-1
so,point of intersection (-2,0) (-1,2)